OpenStudy (anonymous):

[(4^3x^3/4y^4/3)/(16x^-1/2y^2)]^-1/2

OpenStudy (anonymous):

\[\huge{(}\frac{\frac{\frac{4^{3}x^{3}}{4y^{4}}}{3}}{16x^{-\frac{1}{2}}y^{2}})^{-\frac{1}{2}} = (\frac{4^{3}x^{3}}{4y^{4}} \times \frac{3}{16x^{-\frac12}y^{2}})^{-\frac 12}\]\[\huge(\frac{3x^{\frac{7}{2}}}{y^{6}})^{-\frac12} = (\frac{y^{6}}{3x^{\frac{7}{2}}})^{\frac12} = \frac{y^{3}}{3^{\frac12}x^{\frac74}}\]

OpenStudy (anonymous):

@ParthKohli Do you see any errors?

OpenStudy (anonymous):

@katlvb21 Was what I did the original problem?

OpenStudy (anonymous):

thank you for the answer. but it was supposed to be: